# Gradient-Free Adaptive Global Pruning for Pre-trained Language Models

### v1

@@ Abstract
The transformative impact of large language models (LLMs) like LLaMA and GPT on natural language processing is countered by their prohibitive computational demands. Pruning has emerged as a pivotal compression strategy, introducing sparsity to enhance both memory and computational efficiency. Yet, traditional global pruning is impractical for LLMs due to scalability issues, while local pruning, despite its efficiency, leads to suboptimal solutions. Addressing these challenges, we propose Adaptive Global Pruning (AdaGP), a novel framework that redefines the global pruning process into manageable, coordinated subproblems, allowing for resource-efficient optimization with global optimality. AdaGP’s approach, which conceptualizes LLMs as a chain of modular functions and leverages auxiliary variables for problem decomposition, not only facilitates a pragmatic application on LLMs but also demonstrates significant performance improvements, particularly in high-sparsity regimes where it surpasses current state-of-the-art methods. The source code of AdaGP is available at https://github.com/BaiTheBest/AdaGP.
@@ 1 Introduction
Large language models Touvron et al. (2023); OpenAI (2023) have recently transformed the field of natural language processing (NLP) by delivering exceptional results across a variety of intricate language benchmarks Wei et al. (2022); Bommarito II and Katz (2022); Bubeck et al. (2023). Nonetheless, these models, with billions of parameters, generally necessitate significant computational resources. To make large language models (LLMs) more accessible, extensive efforts have been devoted to model compression of LLMs Xu and McAuley (2023); Ling et al. (2023); Bai et al. (2024), including pruning, quantization, knowledge distillation, and low-rank factorization, etc. Pruning, by introducing sparsity, jointly enhances memory and computational efficiency and offers unparalleled flexibility, seamlessly integrating with any LLMs, thus standing out as a highly effective and widely adopted compression strategy.
Figure 1: AdaGP decomposes the global pruning of LLMs into manageable subproblems by leveraging the chain of modules and auxiliary variables.
@@
Model pruning has a long history LeCun et al. (1989) and has proven effective in applications related to vision and smaller language models Hoefler et al. (2021). However, conventional pruning techniques, which rely on global pruning and require loading the entire model into the same GPU Mallya and Lazebnik (2018); Singh and Alistarh (2020), become impractical for today’s LLMs due to their vast size. Recently, several local pruning methods, which consider compressing each layer separately, and the overall compressed model is then obtained by “stitching together” the individually compressed layers, are proposed for billion-scale LLMs. Frantar and Alistarh (2023) developed SparseGPT, an efficient unstructured pruning method for LLMs with hundreds of billions of parameters, achieving up to 60% parameter reduction with minimal performance loss. Sun et al. (2023) introduced a novel pruning criterion in Wanda, which evaluates weights by considering both magnitude and related input activations. However, despite its efficiency gains, local pruning only aims to minimize the local error for each specific layer under sparsity constraints and hence the final solution is suboptimal to the final goal of the global pruning. In other words, beyond the genuine goal of pruning which is just to align the input and output of the pruned and original models, local pruning over-aligns all the intermediate layers’ activations between them, resulting in a suboptimal solution, especially in high-sparsity regimes Singh and Alistarh (2020); Sung et al. (2023).
@@
To address these challenges, namely, to achieve global pruning with low memory consumption, we propose Adaptive Global Pruning (AdaGP), that decomposes the global pruning objective into many subproblems, each of which can be solved using low resources and can coordinate each other toward the global pruning objective. More specifically, we first formulate LLMs as a composite function where the output of one module is the input of the next, based on which we reformulate the global pruning goal into its equivalent form, with auxiliary variables that facilitate its decomposition and its subproblems’ coordination. Then we propose an algorithm that achieves the alternating optimization of the subproblems of all the variables and auxiliary variables with low resource and global optimality, thanks to the closed-form solution of each subproblem. Empirically, we find that AdaGP can consistently improve the performance of local pruning methods, particularly in high sparsity regimes (> 60%), where the perplexity can be significantly decreased by up to around 80% as compared to the state-of-the-art local pruning methods, SparseGPT.
@@
Furthermore, our AdaGP framework can be readily applicable to enhance the performance of most existing local pruning solvers, such as SparseGPT and Wanda, with marginal additional computational overhead. This adaptability ensures that our framework can be seamlessly integrated into a wide range of LLMs and pruning methods, making it a versatile tool and useful baseline for future research exploiting the sparsity of LLMs.
@@ 2 Related Work
Unstructured pruning looks at simplifying the complexity of LLMs by removing certain parameters regardless of the model’s inherent structure. This approach typically involves setting a threshold to nullify parameters below it, leading to a model with a non-uniform sparse structure. Frantar and Alistarh (2023) developed SparseGPT, an efficient unstructured pruning method for LLMs with hundreds of billions of parameters, achieving up to 60% parameter reduction with minimal performance loss. Sun et al. (2023) introduced a novel pruning criterion in Wanda, which evaluates weights by considering both magnitude and related input activations. This approach is beneficial in linear layers of LLMs, helping to identify and remove less significant weights. Tuli and Jha (2023) proposes DynaTran, a dynamic inference scheme for pruning activations at runtime, supported by a specially designed ASIC architecture, AccelTran, to enhance transformer inference throughput.
On the other hand, structured pruning involves the selective removal of groups of weights, where ‘group’ might mean blocks of weights, filters, attention heads, or other structures conducive to hardware acceleration. Ma et al. (2023) introduced the LLM-Pruner, a framework designed for structured pruning of LLMs, which utilizes a combination of first-order data and Hessian information for effective importance estimation. This aids in identifying crucial groups for pruning. Li et al. (2023) proposed LoSparse, a novel approach combining low-rank and sparse matrix approximations to balance pruning and expressive power. Tao et al. (2023) extended this concept to pruning hidden dimensions in LLMs, including embedding layers and attention heads. Kurtic et al. (2023) presented ZipLM, a structured compression method for LLMs, optimizing for compression and accuracy while considering specific hardware constraints.
Our work falls in the category of unstructured pruning of LLMs, where existing methods such as SparseGPT and Wanda only consider an entirely local pruning algorithm and suffer from sub-optimal performance. We discuss the limitations and challenges of entirely local pruning in Sec. 3.
3 Background
@@ 3.1 Global Pruning
Given a pre-trained neural network $f$ and inputs $\mathbf{X}$, global pruning aims to find a global sparsity mask $\mathbf{M}$ and possibly updated weights $\widehat{\mathbf{W}}$ to minimize the global loss between the final outputs of the uncompressed and compressed model:
@@
where $\odot$ denotes the element-wise multiplication. In addition to NP-hardness Blumensath and Davies (2008), however, a critical challenge in solving Eq. 1 for modern billion-scale LLMs is the huge memory requirement which prohibits storing the entire model on a single GPU. Although there exists toolkits, e.g., Paszke et al. (2019); Smith et al. (2022), that enable parallelism across multiple GPUs, such a setup may require resources that are not easily accessible, raising interests in developing other pruning methods that are functional regardless of the availability of computational resources.
@@ 3.2 Local Pruning
Local pruning circumvents the memory issue mentioned above by dividing the full model compression into sub-problems for each layer, $\ell$, and constructing a local loss to measure the $\ell_{2}$-error between the outputs of the uncompressed and compressed layers. Hence, the local pruning can be formulated by
@@
A particularly popular approach is to separate the problem into mask selection and weight reconstruction Hubara et al. (2021); Kwon et al. (2022). Concretely, this means first choosing a pruning mask $\mathbf{M}$ according to some salient criterion, like the weight magnitude Zhu and Gupta (2017), and then optimizing the remaining un-pruned weights while keeping the mask unchanged. Importantly, once the mask is fixed, Eq. 2 turns into a linear regression problem that can be easily optimized.
Figure 2: Illustration of our adaptive-global pruning method compared to conventional global pruning and local pruning. We consider a two-layer neural network as an abstraction for simplicity. Global pruning (left) is memory prohibitive due to poor scalability. Local pruning (mid) considers pruning each layer independently, while inevitably sacrificing performance due to the ignorance of global supervision. Our adaptive global pruning (right) achieves global pruning with low memory cost by leveraging auxiliary variables and soft constraints.
@@
Early work Kingdon and Kingdon (1997) applied iterated linear regression to small networks. Recently, the AdaPrune approach Hubara et al. (2021) has shown good results for this problem on modern models via magnitude-based weight selection, followed by applying SGD steps to reconstruct the remaining weights. Follow-up works demonstrate that pruning accuracy can be further improved by removing the strict separation between mask selection and weight reconstruction. In addition, as we mentioned in Section 2, the work Frantar and Alistarh (2023); Sun et al. (2023) further advance in the scalability and pruning criterion of the pruning process.
3.3 What’s Wrong with Local Pruning?
As shown in Eq. 2, local pruning focuses on minimizing the error for each specific layer $\ell$ subject to sparsity constraints which are embedded in the mask $\mathbf{W}_{\ell}$ and weights $\widehat{\mathbf{W}}_{\ell}$. While the primary goal of the pruning process is to ensure that the input and output of the pruned model align closely with those of the original models, local pruning overly constrains the alignment of input and output of all the intermediate layers between the two models, leading to a performance degradation. This results in a solution that is suboptimal with respect to the global pruning problem.
4 Adaptive Global Pruning for LLMs
We present our proposed method AdaGP that can address the drawbacks of existing pruning methods by achieving a global pruning with low memory consumption. AdaGP decomposes the global pruning into many subproblems, each of which can be solved using low resources and can coordinate each other toward the global pruning objective.
@@ 4.1 Motivation
The development of AdaGP is motivated by the observation: LLMs can be formulated as a composite function such that the output of one module is the input of the next. This allows us to reformulate the global pruning goal into its equivalent form with auxiliary variables that enable the decomposition of the original problem into multiple subproblems, as detailed in Figure 2 and Section 4.2. Then we develop a resource-efficient algorithm that achieves global optimality via alternating optimization for each of the subproblems in which we obtain closed-form solutions for each of the steps. This has been illustrated in Figure 2 and Section 4.3.
@@ 4.2 A Unified Formulation of Pruning
In this section, we present the reformulation of the global pruning problem into an equivalent one by introducing auxiliary variables. This reformulation enables us to swiftly shift between two extremes of global pruning and local pruning and the decomposition of the problem into many independent subproblems.
The key idea behind our reformulation is to decouple the densely parametric parts (linear layers) from the non-parametric parts (activation function, self-attention, and layer norm, etc) using a splitting technique. Formally, given an LLM formulated as a composite function with dense layers $\ell\in[L]:=\{1,2,\ldots,L\}$, where $L$ is the total number of dense layers and define $\Omega\subseteq[L-1]$ to be a subset of layers for which we control the extent of globality. Note that there can be non-parameteric and potentially nonlinear layers between layers $\ell$ and $\ell+1$ in ${L}$ and we denote those layers by function $\phi_{\ell}(\cdot)$. In the reformulation, rather than feeding the output of a dense layer $\ell\in[L]$ directly into a non-parametric and potentially non-linear layer after it, we store the output of layer $\ell$ in a new variable $\mathbf{z}_{\ell}=\mathbf{W}_{\ell}\boldsymbol{a}_{\ell-1}$  [^1] For the sake of simplicity and clearer presentation, the bias term is omitted in the following equations where its exclusion does not lead to confusion., where $\boldsymbol{a}_{\ell-1}$ is the input to layer $\ell$ and can be represented by $\boldsymbol{a}_{\ell-1}=\phi_{\ell-1}(\mathbf{z}_{\ell-1})$. Note that by definition, $\boldsymbol{a}_{0}$ is the input data. Consequently, we can pose the pruning problem as:
@@
where $[L-1]\backslash\Omega$ denotes the complement set of $\Omega$ and $\boldsymbol{a}_{\ell}^{pre}$ and $\mathbf{z}_{\ell}^{pre}$ denote the corresponding values of $\boldsymbol{a}_{\ell}$ and $\mathbf{z}_{\ell}$ obtained by running a forward-propagation from the pre-trained model respectively.
In our proposed reformulation above, the key to understanding its unified nature lies in the interpretation of the set $\Omega$. The versatility of our approach is illustrated in the following remark:
Remark 4.1.
For the given LLM and definition of $\Omega$, our formulation can seamlessly treat both global and local pruning as special cases depending on the choices of $\Omega$. Specifically:
  • When $\Omega=[L-1]$, solving our pruning formulation is equivalent to global pruning, accounting for inter-layer dependencies across the entire network.
  • When $\Omega=\emptyset$, the formulation simplifies to local pruning, considering each layer independently and ignoring inter-layer dependencies.
The ability to shift between these two extremes, and potentially any intermediate configuration, demonstrates the flexibility and comprehensiveness of our reformulation. This flexibility not only caters to a wide range of pruning strategies but also provides a unified framework to compare and contrast the effectiveness of different pruning methods under a consistent mathematical lens.
@@ 4.3 Algorithm Design
In this section, we introduce a choice of $\Omega$ that achieves a good trade-off between global and local pruning and propose an algorithm to solve the resulting problem. Specifically, we propose an adaptive approach where global pruning is applied within each decoder layer, while still adhering to a “local pruning” strategy between different decoder layers. Moreover, given that the feed-forward network (FFN) module in each decoder layer accounts for more than two-thirds of the total parameters Liu et al. (2024), our approach prioritizes the global pruning of these FFN layers. This strategy strikes a balance between the computational feasibility of pruning large-scale models and the effectiveness of the pruning process.
Formally, we select $\Omega$ to be the FFN layers, denoted by $\Omega_{\text{FFN}}$. As a result, we obtain the problem of interest by substituting $\Omega$ with $\Omega_{\text{FFN}}$ in Eq. 3. We have the following problem now:
$\begin{split}&\min_{\{\widehat{\mathbf{{W}}}_{\ell}\},\{\mathbf{M}_{\ell}\},\{\boldsymbol{a}_{\ell}\},\{\mathbf{z}_{\ell}\}}\mathcal{L}(\mathbf{z}_{L},\mathbf{y}),\\ &\text{s.t. \;}\mathbf{z}_{\ell}=(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1},\;\forall\;\ell\in[L],\\ &\quad\;\;\,\boldsymbol{a}_{\ell}=\phi_{\ell}(\mathbf{z}_{\ell}),\;\forall\;\ell\in\Omega_{\text{FFN}},\\ &\quad\;\boldsymbol{a}_{\ell},\mathbf{z}_{\ell}=\boldsymbol{a}_{\ell}^{pre},\mathbf{z}_{\ell}^{pre},\;\forall\;\ell\in[L-1]\backslash\Omega_{\text{FFN}},\end{split}$ (4)
Instead of solving Eq. 4 directly, we relax the constraints by adding an $\ell_{2}$-penalty function to the objective to obtain an unconstrained problem. In addition, by considering the “local pruning" between the decoder layer, we can simplify Eq. 4 to obtain a unconstrained subproblem for each of the decoder layer as:
@@
$\begin{split}&\min_{\{\widehat{\mathbf{{W}}}_{\ell}\},\{\mathbf{M}_{\ell}\},\boldsymbol{a}_{\ell},\mathbf{z}_{\ell}}\alpha\|\mathbf{z}_{\ell+1}^{pre}-(\mathbf{M}_{\ell+1}\odot\widehat{\mathbf{W}}_{\ell+1})\boldsymbol{a}_{\ell}\|_{2}^{2}\\ &\quad\quad\quad+\beta\|\boldsymbol{a}_{\ell}-\phi_{\ell}(\mathbf{z}_{\ell})\|_{2}^{2},\\ &\quad\quad\quad+\alpha\|\mathbf{z}_{\ell}-(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}^{pre}\|_{2}^{2},\end{split}$ (5)
where $\alpha$, $\beta$ are hyperparameters for the penalty of corresponding constraint in Eq. 4 and layers $\ell$ and $\ell+1$ correspond to the up-scaling and down-scaling linear layers in FFN. Note that for the last decoder layer, we replace $\mathbf{z}_{\ell+1}^{pre}$ with $y$.

### v3

# SparseLLM: Towards Global Pruning of Pre-trained Language Models

@@ Abstract
The transformative impact of large language models (LLMs) like LLaMA and GPT on natural language processing is countered by their prohibitive computational demands. Pruning has emerged as a pivotal compression strategy, introducing sparsity to enhance both memory and computational efficiency. Yet, traditional global pruning is impractical for LLMs due to scalability issues, while local pruning, despite its efficiency, leads to suboptimal solutions. Addressing these challenges, we propose SparseLLM, a novel framework that redefines the global pruning process into manageable, coordinated subproblems, allowing for resource-efficient optimization with global optimality. SparseLLM’s approach, which conceptualizes LLMs as a chain of modular functions and leverages auxiliary variables for problem decomposition, not only facilitates a pragmatic application on LLMs but also demonstrates significant performance improvements, particularly in high-sparsity regimes where it surpasses current state-of-the-art methods. Code can be found at https://github.com/BaiTheBest/SparseLLM.
@@ 1 Introduction
Large language models (LLMs) Touvron et al. (2023); OpenAI (2023) have recently transformed the field of natural language processing (NLP) by delivering exceptional results across a variety of intricate language benchmarks Wei et al. (2022); Bommarito II and Katz (2022); Bubeck et al. (2023). Nonetheless, these models, with billions of parameters, generally necessitate significant computational resources. To make LLMs more accessible, extensive efforts have been devoted to model compression of LLMs Xu and McAuley (2023); Bai et al. (2024), including pruning, quantization, knowledge distillation, and low-rank factorization. Pruning, by introducing sparsity, jointly enhances memory and computational efficiency and offers unparalleled flexibility, seamlessly integrating with any LLMs, thus standing out as a highly effective and widely adopted compression strategy.
@@
Model pruning has a long history LeCun et al. (1989) and has proven effective in applications related to vision and smaller language models Hoefler et al. (2021). However, conventional pruning techniques, which rely on global pruning and require loading the entire model into the same GPU Mallya and Lazebnik (2018); Singh and Alistarh (2020), become impractical for today’s LLMs due to their vast size. Recently, several local pruning methods, which consider compressing each layer separately, and the overall compressed model is then obtained by “stitching together” the individually compressed layers, are proposed for billion-scale LLMs. SparseGPT Frantar and Alistarh (2023), an efficient unstructured pruning method for LLMs with hundreds of billions of parameters, achieves up to 60% parameter reduction with minimal performance loss. Another approach, Wanda Sun et al. (2023), introduces a novel pruning criterion that evaluates weights by considering both magnitude and related input activations. Despite its efficiency gains, local pruning only aims to minimize the local error for each specific layer under sparsity constraints, resulting in a suboptimal solultion for the overall model. This is because local pruning over-aligns the intermediate layers’ activations, leading to suboptimal performance, especially in high-sparsity regimes Singh and Alistarh (2020); Sung et al. (2023).
Figure 1: SparseLLM decomposes the global pruning of LLMs into manageable subproblems by leveraging the chain of modules and auxiliary variables while maintaining dependencies.
@@
To address these challenges and achieve global pruning with low memory consumption, we propose SparseLLM that decomposes the global pruning objective into multiple subproblems, each of which can be solved with low resources and coordinate to achieve the global pruning objective. More specifically, we first formulate LLMs as a composite function where the output of one module is the input of the next. Based on this formulation, we reformulate the global pruning goal into an equivalent form with auxiliary variables that facilitate its decomposition and coordination of the subproblems. Then we propose an alternating optimization algorithm to efficiently solve the subproblems, achieving computational resource efficiency and global optimality, due to the close-form solution of each subproblem. Empirically, we find that SparseLLM can consistently improve the performance of local pruning methods, particularly in high sparsity regimes (> 60%), where the perplexity can be significantly decreased by up to around 80% as compared to the state-of-the-art methods.
@@
Furthermore, our SparseLLM framework can be readily applicable to enhance the performance of most existing local pruning solvers, such as SparseGPT and Wanda, with marginal additional computational overhead. This adaptability ensures that our framework can be seamlessly integrated into a wide range of LLMs and pruning methods, making it a versatile tool and useful baseline for future research exploiting the sparsity of LLMs.
@@ 2 Related work
Unstructured pruning looks at simplifying the complexity of LLMs by removing certain parameters regardless of the model’s inherent structure. This approach typically involves setting a threshold to nullify parameters below it, leading to a model with a non-uniform sparse structure. SparseGPT Frantar and Alistarh (2023), an efficient unstructured pruning method for LLMs with hundreds of billions of parameters, achieves up to 60% parameter reduction with minimal performance loss. A novel pruning criterion is introduced in Wanda Sun et al. (2023), which evaluates weights by considering both magnitude and related input activations. This approach is beneficial in linear layers of LLMs, helping to identify and remove less significant weights. Tuli and Jha Tuli and Jha (2023) proposed DynaTran, a dynamic inference scheme for pruning activations at runtime, supported by a specially designed ASIC architecture, AccelTran, to enhance transformer inference throughput.
On the other hand, structured pruning involves the selective removal of groups of weights, where “group” might mean blocks of weights, filters, attention heads, or other structures conducive to hardware acceleration. Ma et al. Ma et al. (2023) introduced the LLM-Pruner, a framework designed for structured pruning of LLMs, which utilizes a combination of first-order data and Hessian information for effective importance estimation. This aids in identifying crucial groups for pruning. Li et al. Li et al. (2023) proposed LoSparse, a novel approach combining low-rank and sparse matrix approximations to balance pruning and expressive power. Tao et al. Tao et al. (2023) extended this concept to pruning hidden dimensions in LLMs, including embedding layers and attention heads. ZipLM Kurtic et al. (2023), a structured compression method for LLMs, is proposed to optimize for compression and accuracy while considering specific hardware constraints.
Our work falls in the category of unstructured pruning of LLMs, where existing methods such as SparseGPT and Wanda only consider an entirely local pruning algorithm and suffer from suboptimal performance. We discuss the limitations and challenges of entirely local pruning in Sec. 3.
3 Background and notation
@@ 3.1 Global pruning
Given a pre-trained neural network $f$ with parameter $\mathbf{W}$ and inputs $\mathbf{X}$, global pruning aims to find a global sparsity mask $\mathbf{M}$ and possibly updated weights $\widehat{\mathbf{W}}$ to minimize the global loss $\mathcal{L}$ between the final outputs of the uncompressed and compressed model:
@@
where $\odot$ denotes the element-wise multiplication. In addition to NP-hardness Blumensath and Davies (2008), however, a critical challenge in solving Eq. 1 is the huge memory cost, as one needs to store the entire model in a single GPU, rendering this method impractical for modern billion-scale LLMs.
@@ 3.2 Local pruning
Local pruning circumvents the memory issue mentioned above by dividing the full model compression into subproblems for each layer and constructing a local loss to measure the $\ell_{2}$-error between the outputs of the uncompressed and compressed layers. Hence, the local pruning can be formulated by
@@
A particularly popular approach is to separate the problem into mask selection and weight reconstruction Hubara et al. (2021); Kwon et al. (2022). Concretely, this means first choosing a pruning mask $\mathbf{M}$ according to some salient criterion, like the weight magnitude Zhu and Gupta (2017), and then optimizing the remaining unpruned weights while keeping the mask unchanged. Importantly, once the mask is fixed, Eq. 2 turns into a linear regression problem that can be easily optimized.
@@
Early work Kingdon and Kingdon (1997) applied iterated linear regression to small networks. Recently, the AdaPrune approach Hubara et al. (2021) has shown good results for this problem on modern models via magnitude-based weight selection, followed by applying SGD steps to reconstruct the remaining weights. Follow-up works demonstrate that pruning accuracy can be further improved by removing the strict separation between mask selection and weight reconstruction. More recently, Frantar and Alistarh (2023) developed SparseGPT, an efficient unstructured pruning method for LLMs with hundreds of billions of parameters, achieving up to 60% parameter reduction with minimal performance loss. Sun et al. (2023) introduced a novel pruning criterion in Wanda, which evaluates weights by considering both magnitude and related input activations.
3.3 What is wrong with local pruning?
As shown in Eq. 2, local pruning focuses on minimizing the error for each specific layer $\ell$ subject to sparsity constraints. This results in a suboptimal solution with respect to the global pruning problem. While the primary goal of pruning is to ensure that the input and output of the pruned model align closely with those of the original models, the local pruning overly constrains the activations of all the intermediate layers between the two models, leading to performance degradation.
4 SparseLLM: Towards global pruning for LLMs
We present our proposed method SparseLLM that can address the drawbacks of existing pruning methods by achieving a global pruning with low memory consumption. SparseLLM decomposes the global pruning objective into many subproblems, each of which can be solved using low resources and can coordinate each other toward the global pruning objective. An overview of SparseLLM on the OPT and LlaMA configurations are shown in Figure 2.
@@ 4.1 Motivation
The development of SparseLLM is motivated by the observation: LLMs can be formulated as a composite function such that the output of one module is the input of the next. This allows us to reformulate the global pruning goal into its equivalent form with auxiliary variables that enable the decomposition into multiple subproblems, as detailed in Sec. 4.2. Then we develop a resource-efficient algorithm that achieves the alternating optimization of the subproblems with global optimality, thanks to the close-form solution of each subproblem, as illustrated in Sec. 4.3.
@@ 4.2 A unified formulation of pruning
In this section, we present the reformulation of the global pruning problem into an equivalent one by introducing auxiliary variables. This reformulation provides a more flexible form and enables the decomposition of the problem into many manageable subproblems.
The key idea behind our formulation is to decouple the densely parametric parts (linear layers) from non-parametric parts (activation function, self-attention, layer norm, etc) using a splitting technique. Rather than feeding the output of the dense linear layer $\mathbf{W}_{\ell}$ directly into the non-parametric and potentially nonlinear layer $\phi_{\ell}$, we store the output of layer $\ell$ in a new variable $\mathbf{z}_{\ell}=\mathbf{W}_{\ell}\boldsymbol{a}_{\ell-1}$  [^1] For the sake of simplicity and clearer presentation, the bias term is omitted in the following equations where its exclusion does not lead to confusion.. We also represent the output of the non-parametric layer as a vector of activations $\boldsymbol{a}_{\ell}=\phi_{\ell}(\mathbf{z}_{\ell})$. We then solve the following problem:
@@
where $L$ represents the total number of dense (linear) layers and $[L]=\{1,2,\cdots,L\}$. $[L-1]\backslash\Omega$ denotes the complement set of $\Omega$. We use $\boldsymbol{a}_{\ell}^{pre}$, $\mathbf{z}_{\ell}^{pre}$ to denote the corresponding intermediate variables’ values of the original dense (i.e., without pruning) pre-trained model. $y$ denotes the ground-truth final output of the dense pre-trained model.
In our proposed formulation above, its unified nature lies in the interpretation and application of the set $\Omega$, which denotes the indices of layers subject to the pruning process. Intuitively, $\Omega$ measures how “global” the pruning is. The bigger the set of $\Omega$ is, the more layers are connected via the second constraint, and the pruning is more towards the global extreme, and vice versa. The generality and versatility of our formulation is illustrated in the following remark:
Remark 4.1 (Generality and flexibility of Eq. 3).
Given an LLM formulated as a composite function with dense layers $l\in\{1,2,\ldots,L-1\}$, where $L$ is the total number of dense layers and $\Omega$ denotes the set of layers subject to the pruning process. Our formulation can seamlessly treat both global and local pruning as special cases under certain conditions. Specifically:
  • When $\Omega=\{1,2,\ldots,L-1\}$, solving our pruning formulation is equivalent to global pruning, accounting for inter-layer dependencies across the entire network.
  • When $\Omega=\emptyset$, the formulation simplifies to local pruning, considering each layer independently (the last constraint dominates and “cuts” all layer dependencies with pre-trained values.)
The ability to shift between these two extremes, and potentially any intermediate configurations, demonstrates the flexibility and comprehensiveness of our formulation. By adjusting $\Omega$, one can seamlessly transition from a global perspective to a local perspective. This flexibility not only caters to a wide range of pruning strategies but also provides a unified framework to compare and contrast the effectiveness of different pruning methods under a consistent mathematical lens.
Figure 2: Illustration of SparseLLM on OPT and LlaMA. The auxiliary variables and soft constraints (i.e., $\approx$) allow SparseLLM to decompose the global pruning into manageable subproblems while maintaining the dependencies. Subproblems are analytically solvable and enjoy fast convergence.
@@ 4.3 Algorithm design
In this section, we introduce the algorithm design of SparseLLM, which alternatively optimizes the subproblems associated with the corresponding variables. This approach is resource-efficient and achieves global optimality, attributed to the closed-form solutions that each subproblem yields.
@@
The key idea of our algorithm lies behind the flexibility of $\Omega$ in our Eq. 3, as we want to find a better trade-off between completely global (memory bottleneck) and completely local (suboptimal performance) pruning. Naively applying SparseLLM to prune all layers globally is impractical. On the other hand, recent work shows that the feed-forward network (FFN) module in each decoder layer accounts for more than two-thirds of the total parameters in an LLM Liu et al. (2024). Therefore, our SparseLLM prioritizes the global pruning of the FFN module, while still adhering to a local pruning strategy for the multi-head attention (MHA) module (see Figure 2). This strategy strikes a balance between the computational feasibility of pruning large-scale models and the effectiveness of the pruning process, adhering to the limitations and practices of state-of-the-art LLM pruning frameworks.
Formally speaking, rather than try to solve Eq. 3 directly, we first relax the constraints by adding an $\ell_{2}$-penalty function to the objective and attack the unconstrained problem:
@@
$\begin{split}\mathcal{L}(\mathbf{z}_{L},\mathbf{y})+\alpha\sum\nolimits_{\ell\in[L]}\|\mathbf{z}_{\ell}-(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}\|_{2}^{2}+\beta\sum\nolimits_{\ell\in\Omega_{\text{FFN}}}\|\boldsymbol{a}_{\ell}-\phi_{\ell}(\mathbf{z}_{\ell})\|_{2}^{2},\end{split}$ (4)
where $\alpha$, $\beta$ are hyperparameters for controlling the weight of each constraint. $\Omega_{\text{FFN}}$ denotes the set of indexes for the linear layers in the FFN module of each decoder layer, i.e., linear layers from the same FFN module are pruned globally. For simplicity, the superscript “pre” of $\boldsymbol{a}_{\ell}$ and $\mathbf{z}_{\ell}$ in the third constraint in Eq. 3 is omitted here, i.e., for $\ell\not\in\Omega_{\text{FFN}}$ the $\boldsymbol{a}_{\ell}$ and $\mathbf{z}_{\ell}$ are fixed and equal to the pre-trained model’s intermediate value in the second term of Eq. 4. In the following subsections, we illustrate how we approach the pruning of FFN and MHA modules, respectively.
4.3.1 SparseLLM on OPT models
For each decoder layer in a pre-trained LLM, our Eq. 4 instantly simplifies to globally pruning the corresponding FFN module within that decoder layer as:
$\begin{split}\alpha\|\mathbf{z}_{\ell+1}^{pre}-(\mathbf{M}_{\ell+1}\odot\widehat{\mathbf{W}}_{\ell+1})\boldsymbol{a}_{\ell}\|_{2}^{2}+\beta\|\boldsymbol{a}_{\ell}-\phi_{\ell}(\mathbf{z}_{\ell})\|_{2}^{2}+\alpha\|\mathbf{z}_{\ell}-(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}^{pre}\|_{2}^{2},\end{split}$ (5)
where layers $\ell$ and $\ell+1$ correspond to the up-projection and down-projection linear layers.